# Semiclassics for the QCD vacuum structure through $T^{2}$-compactification with the baryon-’t Hooft flux

### v1

@@ 3.1 Setup
We compactify the spacetime manifold into $\mathbb{R}^{2}\times T^{2}$ with a minimal ’t Hooft flux for $SU(N)$. As above, we can introduce the minimal ’t Hooft flux by taking the transtion functions $g_{3}(x_{4})$ and $g_{4}(x_{3})$ for $T^{2}$ satisfying
$g_{3}(L)^{\dagger}g_{4}(0)^{\dagger}g_{3}(0)g_{4}(L)=\mathrm{e}^{\frac{2\pi\mathrm{i}}{N}}.$ (11)
Unlike the pure Yang-Mills case, in the presence of a fundamental-representation fermion, the nontrivial ’t Hooft flux gives an inconsistency. To resolve this problem, we need to remember that the symmetry group has the nontrivial quotient with the gauge redundancy,
@@
To see how it works, let us introduce the unit $U(1)_{B}$ flux in the torus $T^{2}$, which is the $1/N$ fractional flux in terms of the quark-number symmetry $U(1)_{V}$ as $U(1)_{B}=U(1)_{V}/\mathbb{Z}_{N}$:
@@ 3.2.1 Low-energy modes
To derive the 2d effective theory, let us determine low-energy modes on small $T^{2}$. For gluons, the boundary conditions are
$a_{\mu}(\vec{x},x_{3}+L,x_{4})=Sa_{1\mu}(\vec{x},x_{3},x_{4})S^{\dagger},~{}~{}~{}~{}a_{\mu}(\vec{x},x_{3},x_{4}+L)=Ca_{1\mu}(\vec{x},x_{3},x_{4})C^{\dagger},$ (18)
@@
whose solutions give perturbatively massless modes. Immediately, we can see the normalizable zero modes should have the positive eigenvalue of $\mathrm{i}\gamma_{3}\gamma_{4}$: $\mathrm{i}\gamma_{3}\gamma_{4}\psi=\psi$, which reduces a 4d spinor to a 2d spinor. After some computations shown in Tanizaki:2022ngt , we obtain one zero-mode per flavor with the boundary condition (20). We thus have $N_{f}$ 2d Dirac fermions as 2d low-energy modes. For later convenience, we bosonize these fermions in the 2d effective theory.
@@ 3.2.2 Center vortices and residual $\mathbb{Z}_{N}$ gauge field
$\mathcal{V}(\vec{x})=``[K\mathrm{e}^{-S_{I}/N}\mathrm{e}^{\mathrm{i}\theta/N}\mathrm{e}^{-\mathrm{i}\varphi/N}]",$ (26)
where $S_{I}$ is the instanton action and $K$ is some appropriate (dimensionful) constant. For $N_{f}\geq 2$, the center-vortex vertex becomes [^3] With the non-Abelian bosonization, the center-vortex vertex should be $\mathcal{V}(\vec{x})=``[K\mathrm{e}^{-S_{I}/N}\mathrm{e}^{i\theta/N}(\operatorname{det}U)^{-1/N}]",$ (27) where $U\in U(N_{f})$ is the non-Abelian bosonized variable. This is nothing but the $1/N$ fractionalization of the Kobayashi-Maskawa-’t Hooft instanton vertex Kobayashi:1970ji ; Kobayashi:1971qz ; Maskawa:1974vs ; tHooft:1976rip .
@@
where we also use the flavor permutation symmetry, a subgroup of $SU(N_{f})$. At first, it appears that they are non-genuine operators as they violate the periodicity of compact bosons, and topological lines attached to them would be required. However, we can avoid this problem by remembering that there is a $\mathbb{Z}_{N}$ gauge field. The key observations are the following two properties:
  • There is $\mathbb{Z}_{N}$ gauge field after the Higgsing by the Polyakov loops. Since the original fermion is in the fundamental representation, the residual $\mathbb{Z}_{N}$ gauge symmetry is the vector-like symmetry $\mathbb{Z}_{N}\subset U(1)_{V}$. Thus, after the bosonization, the coupling should be [^4] Note that this magnetic coupling is consistent with the center-vortex vertex (26). Intuitively, the center vortex introduces a fractional magnetic flux ``$\mathrm{d}a=\frac{2\pi}{N}\delta(\vec{x})"$, which leads to the expression (26) from the magnetic coupling (29).
@@
  • There is a constraint that the topological charge must be an integer. This constraint leads to the sum over the vacuum label $k\in\mathbb{Z}_{N}$, as seen in (9).
These two properties leads to the extension of $\varphi$ (or $\varphi_{1}+\cdots+\varphi_{N_{f}}$) when integrating out gluon’s center contributions, and we denote the extended $2\pi N$-periodic fields as $\eta^{\prime}$: [^5] The extension happens also when we perform the non-Abelian bosonization. In that case, the bosonized variable is a $U(N_{f})$-valued field, which can be expressed as $(\varphi,\tilde{U})\in\mathbb{R}\times SU(N_{f})$ with the identification $(\varphi,\tilde{U})\sim(\varphi+2\pi,\mathrm{e}^{-\frac{2\pi\mathrm{i}}{N_{f}}}\tilde{U})$. The phase of $\operatorname{det}U$ is given by $\varphi$. After the $\mathbb{Z}_{N}$ gauging, the phase of $\operatorname{det}U$ becomes $2\pi N$-periodic $\eta^{\prime}$. Then, the identification is modified as $(\eta^{\prime},\tilde{U})\sim(\eta^{\prime}+2\pi N,\mathrm{e}^{-\frac{2\pi N\mathrm{i}}{N_{f}}}\tilde{U})$, and the chiral effective field is no longer $U(N_{f})$-valued.
@@
    $V[\varphi]=-m\mu\cos(\eta^{\prime})-2K\mathrm{e}^{-S_{I}/N}\cos\left(\frac{\eta^{\prime}-\theta}{N}\right).$ (34)
@@ 4 Lesson about the chiral Lagrangian with $\eta^{\prime}$ in $4$d
The non-Abelian-bosonized 2d effective theory is strikingly similar to the chiral Lagrangian as observed in Ref. Tanizaki:2022ngt (despite the fact that the chiral symmetry cannot be spontaneously broken in 2d due to Coleman-Mermin-Wagner theorem666 This subtlety reflects that the theory $\mathbb{R}^{4}$ is not always equivalent to the theory on $\mathbb{R}^{2}\times T^{2}$ at large $T^{2}$ in the naive way when the system is gapless. We have to take into account how the IR description is modified when we compactify $\mathbb{R}^{4}$ to $\mathbb{R}^{2}\times T^{2}$ with a large-but-finite torus size. In our case, as we well see in Section 4.1, the chiral Lagrangian on $\mathbb{R}^{2}\times T^{2}$ gives the $SU(N_{f})_{1}$ WZW theory, so the adiabatic continuity between the theories at large and small $T^{2}$ appears to work well. ), and it is consistent with the $T^{2}$ flux compactification of the chiral Lagrangian (Sec. 4.1). With this identification, the extended $2\pi N$-periodic compact boson corresponds to the $\eta^{\prime}$ meson. Then, we see that the extension of periodicity is important to couple $\eta^{\prime}$ to the $U(1)_{B}$ background also in $4$ dimensions (Sec. 4.2), and we propose the $4$d chiral Lagrangian with extended $\eta^{\prime}$ that can reproduce the discrete baryon-color-flavor anomaly of Refs. Gaiotto:2017tne ; Tanizaki:2017mtm ; Tanizaki:2018wtg ; Anber:2019nze .
@@ 4.2 Extended periodicity, $U(1)_{B}$ background, and discrete anomaly
In the first place, we determine the coupling of the $U(1)_{B}$ background to the chiral field with $\eta^{\prime}$. To this end, we first see the coupling between the $U(1)_{V}$ background $A_{U(1)}~{}(=A_{B}/N)$ and a $U(N_{f})$ chiral field $U$. We can determine this coupling by observing the WZW term, which gives, in the 4d notation,
$\Gamma_{WZW}$ $\supset\frac{\mathrm{i}N}{24\pi^{2}}\int_{M_{4}}\left\{A_{U(1)}\wedge\operatorname{tr}\left(U^{-1}\mathrm{d}U\right)^{3}+3~{}\mathrm{d}A_{U(1)}\wedge\mathrm{d}A_{U(1)}\wedge\operatorname{tr}\left(U^{-1}\mathrm{d}U\right)\right\}$
$=\frac{\mathrm{i}}{24\pi^{2}}\int_{M_{4}}\left\{A_{B}\wedge\operatorname{tr}\left(U^{-1}\mathrm{d}U\right)^{3}+\frac{3}{N}~{}\mathrm{d}A_{B}\wedge\mathrm{d}A_{B}\wedge\mathrm{d}(\ln\operatorname{det}U)\right\},$ (38)
where the trace is taken over the flavor indices. This can also be obtained by superficially employing the Goldstone-Wilzeck baryon current (with the charge matrix $Q=\frac{1}{N}\textbf{1}_{N_{f}\times N_{f}}$) Witten:1983tw ; Goldstone:1981kk . Here, the first term is the well-known Skyrmion current, which only couples to the $SU(N_{f})$ part. The second term describes the coupling of $\eta^{\prime}(=\ln\operatorname{det}U)$ to the baryon background. The coupling to the baryon instanton number is natural because $\eta^{\prime}$ is the pseudo-Nambu-Goldstone boson of anomalously-broken $U(1)_{\mathrm{axial}}$, which has an extra anomaly in the presence of the $U(1)_{V}$ background [^7] This $1/N$ fractional coupling describes the fractional quantum Hall effect on the domain walls Komargodski:2018odf . Moreover, upon the $T^{2}$ compactification with the nontrivial baryon flux, this coupling reproduces (29) of the 2d effective theory, $\frac{\mathrm{i}}{8\pi^{2}N}\int_{\mathbb{R}^{2}\times T^{2}}\eta^{\prime}\mathrm{d}A_{B}\wedge\mathrm{d}A_{B}\Rightarrow\frac{\mathrm{i}}{2\pi N}\int_{\mathbb{R}^{2}}\eta^{\prime}\mathrm{d}A_{B}=\frac{\mathrm{i}}{2\pi}\int_{\mathbb{R}^{2}}\varphi\,\mathrm{d}A_{U(1)}.$ However, this coupling to $\eta^{\prime}$,

### v3

@@ 3.1 Setup
We compactify the spacetime manifold into $\mathbb{R}^{2}\times T^{2}$ with a minimal ’t Hooft flux for $SU(N)$. As above, we can introduce the minimal ’t Hooft flux by taking the transition functions $g_{3}(x_{4})$ and $g_{4}(x_{3})$ for $T^{2}$ satisfying
$g_{3}(L)^{\dagger}g_{4}(0)^{\dagger}g_{3}(0)g_{4}(L)=\mathrm{e}^{\frac{2\pi\mathrm{i}}{N}}1_{N\times N}.$ (11)
Unlike the pure Yang-Mills case, in the presence of a fundamental-representation fermion, the nontrivial ’t Hooft flux gives an inconsistency. Indeed, for a fundamental quark $\psi(x_{3},x_{4})$, there are two ways to connect $\psi(0,0)$ to $\psi(L,L)$ using transition functions: $\psi(0,0)=g_{3}(0)g_{4}(L)\psi(L,L)=g_{4}(0)g_{3}(L)\psi(L,L)$, which implies $g_{3}(L)^{\dagger}g_{4}(0)^{\dagger}g_{3}(0)g_{4}(L)=1_{N\times N}$. We cannot introduce the ’t Hooft flux alone in this case. To resolve this problem, we need to remember that the symmetry group has the nontrivial quotient with the gauge redundancy,
@@
To see how it works, let us introduce the unit flux of the baryon-number symmetry $U(1)_{B}$ in the torus $T^{2}$, which is the $1/N$ fractional flux in terms of the quark-number symmetry $U(1)_{V}$ as $U(1)_{B}=U(1)_{V}/\mathbb{Z}_{N}$:
@@ 3.2.1 Low-energy modes
To derive the 2d effective theory, let us determine low-energy modes on small $T^{2}$. For the gauge field, the boundary conditions are
@@
whose solutions give perturbatively massless modes. Immediately, we can see the normalizable zero modes should have the positive eigenvalue of $\mathrm{i}\gamma_{3}\gamma_{4}$: $\mathrm{i}\gamma_{3}\gamma_{4}\psi=\psi$, which reduces a 4d spinor to a 2d spinor. After some computations shown in Tanizaki:2022ngt , we obtain one zero-mode per flavor with the boundary condition (20). One can also determine the number of zero-modes by applying the index theorem on $T^{2}$. We thus have $N_{f}$ 2d Dirac fermions as 2d low-energy modes. For later convenience, we bosonize these fermions in the 2d effective theory.
@@ 3.2.2 Center vortices and residual $\mathbb{Z}_{N}$ gauge field
$\mathcal{V}(\vec{x})=``[K\mathrm{e}^{-S_{I}/N}\mathrm{e}^{\mathrm{i}\theta/N}\mathrm{e}^{-\mathrm{i}\varphi/N}],"$ (26)
where $S_{I}$ is the instanton action and $K$ is some appropriate (dimensionful) constant. We put the double quotations to indicate that careful consideration is required to define the fractional vertex $\mathrm{e}^{-\mathrm{i}\varphi/N}$, which is discussed shortly thereafter. For $N_{f}\geq 2$, the center-vortex vertex becomes [^7] With the non-Abelian bosonization, the center-vortex vertex should be $\mathcal{V}(\vec{x})=``[K\mathrm{e}^{-S_{I}/N}\mathrm{e}^{i\theta/N}(\operatorname{det}U)^{-1/N}],"$ (27) where $U\in U(N_{f})$ is the non-Abelian bosonized variable. This is nothing but the $1/N$ fractionalization of the Kobayashi-Maskawa-’t Hooft instanton vertex Kobayashi:1970ji ; Kobayashi:1971qz ; Maskawa:1974vs ; tHooft:1976rip .
@@
where we also use the flavor permutation symmetry, a subgroup of $SU(N_{f})$. At first, it appears that they are non-genuine operators as they violate the periodicity of compact bosons, and topological lines attached to them would be required. However, we can resolve this problem by remembering that there is a $\mathbb{Z}_{N}$ gauge field. The key observations are the following two properties:
  • There is $\mathbb{Z}_{N}$ gauge field after the Higgsing by the Polyakov loops. Since the original fermion is in the fundamental representation, the residual $\mathbb{Z}_{N}$ gauge symmetry is the vector-like symmetry $\mathbb{Z}_{N}\subset U(1)_{V}$. In the bosonization dictionary, the fermion vector-like symmetry is mapped into the magnetic (winding) symmetry of the bosonized variable. The fermion number is translated as the number of kinks, which is the winding number. Thus, after the bosonization, the coupling should be [^8] Note that this magnetic coupling is consistent with the center-vortex vertex (26). Intuitively, the center vortex introduces a fractional magnetic flux ``$\mathrm{d}a=\frac{2\pi}{N}\delta(\vec{x})"$, which leads to the expression (26) from the magnetic coupling (29).
@@
  • There is a constraint that the total topological charge must be an integer. This constraint leads to the sum over the vacuum label $k\in\mathbb{Z}_{N}$, as seen in (9).
These two properties leads to the extension of $\varphi$ (or $\varphi_{1}+\cdots+\varphi_{N_{f}}$) when integrating out gluon’s center contributions, and we denote the extended $2\pi N$-periodic fields as $\eta^{\prime}$: [^9] The extension happens also when we perform the non-Abelian bosonization. In that case, the bosonized variable is a $U(N_{f})$-valued field, which can be expressed as $(\varphi,\tilde{U})\in\mathbb{R}\times SU(N_{f})$ with the identification $(\varphi,\tilde{U})\sim(\varphi+2\pi,\mathrm{e}^{-\frac{2\pi\mathrm{i}}{N_{f}}}\tilde{U})$. The phase of $\operatorname{det}U$ is given by $\varphi$ as $(\varphi,\tilde{U})\mapsto U=\mathrm{e}^{\mathrm{i}\varphi/N_{f}}\tilde{U}\in U(N_{f})$. After the $\mathbb{Z}_{N}$ gauging, the phase of $\operatorname{det}U$ becomes $2\pi N$-periodic $\eta^{\prime}$. Then, the identification is modified as $(\eta^{\prime},\tilde{U})\sim(\eta^{\prime}+2\pi N,\mathrm{e}^{-\frac{2\pi N\mathrm{i}}{N_{f}}}\tilde{U})$, and the chiral effective field is no longer $U(N_{f})$-valued. , [^10] This extension says that the $2\pi N$-winding vortex is minimal as the genuinely codim-$2$ object, but we would like to note that the $2\pi$-winding vortex also exists as a boundary of one-higher-dimensional object. Indeed, the $2\pi$-winding $\eta^{\prime}$ vortex violates the $\mathbb{Z}_{N}$ gauge invariance due to the magnetic coupling (29), so it has to be an endpoint of the open $\mathbb{Z}_{N}$ Wilson loop to be gauge invariant. This is the $2$d counterpart of the “pancake” baryon proposed in Ref. Komargodski:2018odf , where the $2\pi$-winding $\eta^{\prime}$ vortex lives at the edge of the Chern-Simons-enriched domain wall. More intuitively, we may say that $\eta^{\prime}$ extends its periodicity by “eating” the $\mathbb{Z}_{N}$ label for the Yang-Mills confining vacua, so it must jump by $2\pi$ when the original vacuum label jumps at some dislocation. This would clarify why the “pancake” is needed in the proposal of Komargodski:2018odf .
@@
To be concrete, let us explicitly see how the $2$d effective theory is derived for $N_{f}=1$ case. As in (9), the dilute gas approximation with residual $\mathbb{Z}_{N}$ gauging yields,
$Z$ $=\int{\mathcal{D}}a_{\mathbb{Z}_{N}}\int\mathcal{D}\varphi\sum_{k\in\mathbb{Z}_{N}}\mathrm{e}^{\frac{\mathrm{i}}{2\pi}\int a_{\mathbb{Z}_{N}}\wedge\mathrm{d}\varphi}$
×$~{}~{}~{}~{}~{}\times\exp\left[-S[\varphi]+\int d^{2}\vec{x}\left\{2K\mathrm{e}^{-S_{I}/N}\cos\left(\frac{\varphi(\vec{x})-2\pi k}{N}\right)\right\}\right]$ (34)
As explained above, the integration over the residual gauge $a_{\mathbb{Z}_{N}}$ constrains the winding number of $\varphi$: $\int\mathrm{d}\varphi\in 2\pi N\mathbb{Z}$, which allows us to lift $\varphi$ to a $2\pi N$-periodic variable. Here, there is $\mathbb{Z}_{N}$ ambiguity to lift $\varphi$: for given $\varphi$, we have $N$ corresponding $\mathbb{R}/2\pi N\mathbb{Z}$ variable $\varphi\rightarrow\eta^{\prime}=\varphi-2\pi k~{}~{}(k\in\mathbb{Z}_{N})$. Based on this 1-to-$N$ correspondence, we notice the identity:
$\int_{\eta^{\prime}\sim\eta^{\prime}+2\pi N}\mathcal{D}\eta^{\prime}~{}(\cdots)=\int_{\left[\frac{\mathrm{d}\varphi}{2\pi}\right]\in NH^{1}(M_{2};\mathbb{Z})}\mathcal{D}\varphi\sum_{k\in\mathbb{Z}_{N}}\left.(\cdots)\right|_{\eta^{\prime}=\varphi-2\pi k},$ (35)
where $\left[\frac{\mathrm{d}\varphi}{2\pi}\right]\in NH^{1}(M_{2};\mathbb{Z})$ means the constraint that the winding number of $\varphi$ is a multiple of $N$. This identity enables us to rewrite the partition function neatly:
$\int{\mathcal{D}}a_{\mathbb{Z}_{N}}\int\mathcal{D}\varphi\sum_{k\in\mathbb{Z}_{N}}\mathrm{e}^{\frac{\mathrm{i}}{2\pi}\int a_{\mathbb{Z}_{N}}\wedge\mathrm{d}\varphi}(\cdots)\Rightarrow\int_{\eta^{\prime}\sim\eta^{\prime}+2\pi N}\mathcal{D}\eta^{\prime}\left.(\cdots)\right|_{\varphi-2\pi k\mapsto\eta^{\prime}}.$ (36)
@@
    $V[\eta^{\prime}]=-m\mu\cos(\eta^{\prime})-2K\mathrm{e}^{-S_{I}/N}\cos\left(\frac{\eta^{\prime}-\theta}{N}\right).$ (37)
@@ 4 Lesson about the chiral Lagrangian with $\eta^{\prime}$ in $4$d
The non-Abelian-bosonized 2d effective theory is strikingly similar to the chiral Lagrangian as observed in Ref. Tanizaki:2022ngt (despite the fact that the chiral symmetry cannot be spontaneously broken in 2d due to Coleman-Mermin-Wagner theorem111111 This subtlety reflects that the theory on $\mathbb{R}^{4}$ is not always equivalent to the theory on $\mathbb{R}^{2}\times T^{2}$ at large $T^{2}$ in the naive way when the system is gapless. We have to take into account how the IR description is modified when we compactify $\mathbb{R}^{4}$ to $\mathbb{R}^{2}\times T^{2}$ with a large-but-finite torus size. In our case, as we well see in Section 4.1, the chiral Lagrangian on $\mathbb{R}^{2}\times T^{2}$ gives the $SU(N_{f})_{1}$ WZW theory, so the adiabatic continuity between the theories at large and small $T^{2}$ appears to work well. ), and it is consistent with the $T^{2}$ flux compactification of the chiral Lagrangian (Sec. 4.1)121212 We note that the consistency holds both for $N\geq 3$ and $N=2$. Even for the two-color case ($N=2$), the Pauli-Gürsey extended $SU(2N_{f})$ symmetry reduces to the usual chiral symmetry $SU(N_{f})_{L}\times SU(N_{f})_{R}\times U(1)_{V}$ as we introduce the $U(1)_{B}$ flux $A_{B}$ along the $T^{2}$ direction. . With this identification, the extended $2\pi N$-periodic compact boson corresponds to the $\eta^{\prime}$ meson. Then, we see that the extension of periodicity is important to couple $\eta^{\prime}$ to the $U(1)_{B}$ background also in $4$ dimensions (Sec. 4.2), and we propose the $4$d chiral Lagrangian with extended $\eta^{\prime}$ that can reproduce the discrete baryon-color-flavor anomaly of Refs. Gaiotto:2017tne ; Tanizaki:2017mtm ; Tanizaki:2018wtg ; Anber:2019nze .
@@ 4.2 Extended periodicity, $U(1)_{B}$ background, and discrete anomaly
In the first place, we determine the coupling of the $U(1)_{B}$ background to the chiral field with $\eta^{\prime}$. To this end, we first see the coupling between the $U(1)_{V}$ background $A_{U(1)}~{}(=A_{B}/N)$ and a $U(N_{f})$ chiral field $U$. We can determine this coupling by observing the WZW term, which gives, in the 5d notation,
$\Gamma_{WZW}$ $\supset\frac{\mathrm{i}N}{24\pi^{2}}\int_{M_{5}}\left\{\mathrm{d}A_{U(1)}\wedge\operatorname{tr}\left(U^{-1}\mathrm{d}U\right)^{3}+3~{}\mathrm{d}A_{U(1)}\wedge\mathrm{d}A_{U(1)}\wedge\operatorname{tr}\left(U^{-1}\mathrm{d}U\right)\right\}$
$=\frac{\mathrm{i}}{24\pi^{2}}\int_{M_{5}}\left\{\mathrm{d}A_{B}\wedge\operatorname{tr}\left(U^{-1}\mathrm{d}U\right)^{3}+\frac{3}{N}~{}\mathrm{d}A_{B}\wedge\mathrm{d}A_{B}\wedge\mathrm{d}(\ln\operatorname{det}U)\right\},$ (41)
where the trace is taken over the flavor indices. This can also be obtained by superficially employing the Goldstone-Wilzeck baryon current (with the charge matrix $Q=\frac{1}{N}\textbf{1}_{N_{f}\times N_{f}}$) Witten:1983tw ; Goldstone:1981kk . Here, the first term is the well-known Skyrmion current, which only couples to the $SU(N_{f})$ part. The second term describes the coupling of $\eta^{\prime}(=\ln\operatorname{det}U)$ to the baryon background. The coupling to the baryon instanton number is natural because $\eta^{\prime}$ is the pseudo-Nambu-Goldstone boson of anomalously-broken $U(1)_{\mathrm{axial}}$, which has an extra anomaly in the presence of the $U(1)_{V}$ background [^13] This $1/N$ fractional coupling describes the fractional quantum Hall effect on the domain walls Komargodski:2018odf (see also footnote 10). Moreover, upon the $T^{2}$ compactification with the nontrivial baryon flux, this coupling reproduces (29) of the 2d effective theory, $\frac{\mathrm{i}}{8\pi^{2}N}\int_{\mathbb{R}^{2}\times T^{2}}\eta^{\prime}\mathrm{d}A_{B}\wedge\mathrm{d}A_{B}\Rightarrow\frac{\mathrm{i}}{2\pi N}\int_{\mathbb{R}^{2}}\eta^{\prime}\mathrm{d}A_{B}=\frac{\mathrm{i}}{2\pi}\int_{\mathbb{R}^{2}}\varphi\,\mathrm{d}A_{U(1)}.$ However, this coupling to $\eta^{\prime}$,